### v1

# Skull-to-Face: Anatomy-Guided 3D Facial Reconstruction and Editing

Abstract
Deducing the 3D face from a skull is an essential but challenging task in forensic science and archaeology. Existing methods for automated facial reconstruction yield inaccurate results, suffering from the non-determinative nature of the problem that a skull with a sparse set of tissue depth cannot fully determine the skinned face. Additionally, their texture-less results require further post-processing stages to achieve a photo-realistic appearance. This paper proposes an end-to-end 3D face reconstruction and exploration tool, providing textured 3D faces for reference. With the help of state-of-the-art text-to-image diffusion models and image-based facial reconstruction techniques, we generate an initial reference 3D face, whose biological profile aligns with the given skull. We then adapt these initial faces to meet the statistical expectations of extruded anatomical landmarks on the skull through an optimization process. The joint statistical distribution of tissue depths is learned on a small set of anatomical landmarks on the skull. To support further adjustment, we propose an efficient face adaptation tool to assist users in tuning tissue depths, either globally or at local regions, while observing plausible visual feedback. Experiments conducted on a real skull-face dataset demonstrated the effectiveness of our proposed pipeline in terms of reconstruction accuracy, diversity, and stability. Our project page is https://xmlyqing00.github.io/skull-to-face-page/.
Keywords:
Skull-to-Face Reconstruction Human Face Tissue Depth Modeling Human Face Adaptation
1 Introduction
3D facial reconstruction from a skull is an effective approach used by forensic investigators to identify skeletal remains when other information is unavailable. It has been successfully applied in forensic cases [52] and anthropological studies [1]. Given a 3D skull scan, the objective is to derive a 3D face that aligns with both the geometric features and the biological profile of the skull. The anatomic setting of the reconstruction process is illustrated in Fig. 1: A set of predefined skull landmarks [11] is distributed on the query skull. The tissue depths (blue sticks) on landmarks are estimated using anatomical statistic data. The endpoints of these tissue depths, i.e., the facial landmarks (green spheres), depict a coarse geometry shape of the 3D face. A plausible 3D face geometry should be created, satisfying the constraints of these facial landmarks.
Figure 1: Given an input skull scan, based on skull landmarks and tissue depth (blue sticks), facial landmarks (green spheres) can be obtained. The 3D face is inferred under the constraints of facial landmarks.
Traditional facial reconstruction in forensic applications is performed manually using plasticine or clay [53, 25]. Recent approaches [5, 35, 55] employ a digital environment to replace these manual clay modeling methods. Both physical and digital methods for facial reconstruction require years of anatomy and artistic modeling experience, and the results are often unreliable due to the subjective factors [18, 56]. Besides the geometry, a vivid facial texture may be inferred from (e.g., ancestry, age of) the skull. Such realistic facial textures can be desirable in anthropological and forensic applications, but their creation often requires further post-processing steps and significant extra effort.
In the past two decades, researchers have been studying computer-aided skull-to-face reconstruction algorithms to produce 3D faces. They access the problem from various perspectives and multidisciplinary backgrounds, from data-driven [10, 39, 8, 12, 16, 56] to physical simulation [24, 26, 36, 11], providing solid research baselines. Nevertheless, (1) all these face generation methods overlooked multi-modal characteristics beyond the geometry of the skull, such as gender, age, and ancestry. (2) Without modeling the distribution of tissue depth, these methods addressed the problem by solving a deterministic mapping from one skull to one face with statistically average tissue depths [43, 44, 48, 49], making it hard for further tuning in real scenarios.
Our proposed work builds upon two main observations. (1) Firstly, the biological profile, encompassing ancestry, age, gender, and stature of the body, can be estimated from skulls [47]. Different population groups exhibit distinct facial priors, and we argue that incorporating these is crucial for accurate facial reconstruction. We generate an initial guess of the 3D face with these attributes, whereas existing methods often overlook this key information. (2) Secondly, the combinations of tissue depths in real human faces follow intrinsic distributions that have not been well-studied. We introduce an additional control based on the distribution of tissue depth over facial reconstruction, providing an intuitive and effective tool for 3D face manipulation and exploration.
Our approach includes an offline tissue depth distribution training stage and two online inference stages: semantic face model initialization, and anatomy-guided face adaptation. We also develop landmark-based face editing as an optional adjustment stage. (1) During the preprocessing stage, tissue depth distribution can be analyzed using available datasets of scanned skull and face pairs, or obtained from statistical data in forensic anthropology literature [43, 44, 48, 49]. (2) When processing a query skull, the key biological profile of the skull is used to generate a 2D realistic facial image, which is then reconstructed into a textured 3D face. Then, facial landmarks are generated, guided by the pre-trained tissue depth distribution and the geometry of the query skull. Finally, the face geometry is adjusted to align with these facial landmarks, resulting in the final skull-conformed face. (3) As an optional step, we provide a tool that allows users to intuitively adjust the face geometry by tuning a one-dimensional feature in the distribution space of tissue depths.
The main contributions of this work are as follows: (1) We present a first approach that incorporates the biological profile into skull-to-face reconstruction. (2) By analyzing the distribution of tissue depth, we devise an additional control mechanism over the automatic reconstruction workflow, offering an intuitive and effective way to tune the face geometry. (3) Compared to existing methods, our method achieves state-of-the-art geometric accuracy qualitatively and quantitatively. We will release the codes and pre-trained models, making it the first publicly accessible approach in this field.
2 Related Work
2.1 Facial Reconstruction from Skull
Data-driven facial reconstruction.
Statistics-based methods for facial reconstruction employ high-dimensional vectors, typically representing faces through their 3D coordinates or feature points. These methods, using techniques such as PCA, learn the mapping between corresponding skulls and faces from training datasets [10, 39, 8, 12, 57, 33, 16, 22]. A new face is then generated as a linear combination of principal components. Other approaches [33, 14, 6, 39, 50, 30, 40] involve regression algorithms to predict faces from skulls. Zhang et al. [56] project 3D skulls to 2D images and leverage image-to-image style transferring networks [20, 58] for facial reconstruction. However, these data-driven methods require a significant number of joint skull-face scans, which are costly and difficult to acquire. The limited available datasets often result in overfitted networks and dataset bias.
Deformation-based facial reconstruction.
Another category of reconstruction algorithms employs 3D modeling techniques to digitally mimic the manual reconstruction process. A prevalent strategy involves deforming a template face surface [11, 36] or tissue volume [51, 26, 24] to fit with the subject’s skull. This deformation is controlled by imposing penalties on certain geometric smoothness energy to regularize the transformation while adhering to the positional constraints of facial landmarks.
2.2 Digital Face Synthesis
Constrained face editing.
Constrained face editing or re-synthesis techniques involve generating new faces that are both realistic and adhere to specific constraints. Many recent 2D-based methods [2, 27, 38, 13, 46, 45, 19] accomplish this by manipulating semantic codes in the latent space to generate new 2D face images. However, these 2D methods lack underlying 3D face geometry, making it difficult to enforce fine geometry constraints during face editing.
Large face generation models.
Recent years have witnessed significant advancements in large face generation models, including text-to-image generation models (e.g. Stable Diffusion XL [4] and DALL-E2 [41]), and 2D image to 3D face generation models (e.g., GaGAN [28], CoMA [42], DECA [15], EG3D [9], and PanoHead [3]). A recent study, Rodin [54], proposed a text-to-3D-face pipeline, but the synthesized faces are somewhat cartoon-like and unrealistic. NeRF-based methods such as EG3D and PanoHead are less effective in editing 3D geometry (to meet specific geometric constraints).
3 Method
Figure 2: An overview of the proposed Skull-to-Face pipeline, which includes three modules. The Face Generation module synthesizes an initial 3D face in consideration of the biological profile of the skull such as age, gender, and ancestry. The Tissue Depth Distribution (TDD) can suggest a valid combination of facial landmarks as geometry constraints. The Anatomy-guided Face adaptation module optimizes the initial face as the facial reconstruction according to the landmark constraints.
3.1 Pipeline Overview
Our skull-to-face generation pipeline has three components, as is shown in Fig.2. In the Tissue Depth Distribution (TDD) modeling, we analyze the tissue depth distribution in the training set and use the distribution to suggest a valid combination of tissue depths for the skull (Sec. 3.2). Then, when reconstructing a face from a query skull, we have two stages. In the Initial Face Generation stage, we utilize key biological profiles about the human body, i.e., gender, age, and ancestry to generate a 2D candidate portrait using Stable Diffusion XL [4] and lift it to a 3D face using the FLAME model [15] (Sec. 3.3). Finally, with the facial landmarks computed from the tissue depth model as the geometry constraints, we propose an Anatomy-guided Face Adaptation method to modify the 3D face according to facial landmark constraints. The details of the reconstructed face are preserved from the 2D candidate portrait image (Sec. 3.4).
3.2 Distribution of Tissue Depth
Tissue depths define the displacement between facial landmarks and their corresponding skull landmarks. The combination of tissue depths controls the shape of the reconstructed face, which is critical for the fidelity of facial reconstruction and directly influences the reconstruction error. Instead of directly learning the mapping from skull to face as other works [22, 56] did, we analyzed and explored the intrinsic pattern of the tissue depth distribution. We then used the distribution pattern to guide facial reconstruction. This also provides the benefits for users to further adjust tissue depths to explore various plausible faces on skulls.
On each skull $J$, a set of landmarks $\mathcal{P}^{s}=\{\mathbf{p}^{s}_{i}\in\mathbb{R}^{2},i=1,\ldots,n\}$ are defined and can be extracted. In this work, we follow [11] to extract $n=78$ landmarks. Each such skull landmark $\mathbf{p}^{s}_{i}$ is extended outward along the normal directions $\hat{\mathbf{n}}^{s}_{i}$ with a tissue depth length $d_{i}$. The extended landmarks are referred to as facial landmarks $\mathcal{P}^{f}=\{\mathbf{p}^{f}_{i}|\mathbf{p}^{f}_{i}=\mathbf{p}^{s}_{i}+d_{i}\hat{\mathbf{n}}^{s}_{i}\}$, which provide a sparse constraint of the face shape. $\mathcal{D}=\{d_{i}\}$ represents the tissue depths of a person, and $\mathcal{X}=\{\mathcal{D}_{m}\}\in\mathbb{R}^{n}$ represents the collection of tissue depths of $m$ people in the training dataset.
(a) Three representative values of the first principal component are visualized overlaid on the tissue depth distribution.
(b) Adjusting the values of the first principal component provides easy control of tissue depths, and the face shapes change from thin to fat.
Figure 3: The first principal component can meaningfully provide control on the thickness of tissue depths. (a) shows the tissue depth distribution $C$ in the first principal component. Vertical lines represent three representative values, signifying three different combinations of tissue depths, resulting in three types of face shapes: thin, normal, and fat, as illustrated in (b).
3.2.1 TDD-Global Model
We analyze the tissue depth distribution using the training set of the skull-face dataset, which comprises $m$ samples, each with $n=78$ landmarks. It is important to note that the number of training samples $m$ in the publicly accessible skull-face dataset is limited ($m\leq 100$ in our experiments). Consequently, employing a categorized or neural-based model to fit the landmark distribution may be inappropriate. We therefore designed a simple yet effective way to model and control the tissue depths.
We first used principle component analysis (PCA) to project the joint distribution of tissue depths from an $n-$dimension space to a low-dimensional space. We observed that the first principal component predominantly governs the tissue depth distribution, accounting for $51.9\%$ of the variance. This finding significantly reduces the degree of freedom for the combination of 78 tissue depths. We projected joint tissue depths to the first component axis, denoted as $C$. The distribution of $C$ is visualized in Fig. 3(a). Furthermore, the effect of adjusting the values of the first principal component is illustrated In Fig. 3(b), where the directions of the tissues are depicted as blue sticks and the tissue depths in the training set are green spheres. A more comprehensive discussion of PCA components is included in the Supplementary.
When we monotonously sample values from the distribution $C$, the tissue depths change from short to long, and the corresponding face shape changes from thinner to rounder. To better visualize the relationship between face shape and the distribution $C$, we select mean values from the top, middle, and bottom $33\%$ of $C$ as representative combinations of tissue depths, and visualize them in Fig. 3(b). Note that while some related morphable face models also provide correlations between PCA components and face fullness [7], our model is the first model explicitly guided by anatomy, connecting facial skin to the underlying skull.
The above model enables us to modify face shapes based on the joint distributions of all the tissue depths on a skull. And we denote it as the TDD-Global model. Specifically, (a) we can use $C=0$ to sample the average tissue depths when we would like to generate average faces with moderate facial thickness. (b) we also can vary values in the first principal component to sample multiple plausible tissue depth combinations, thereby exploring different face shapes. The effectiveness of TDD-Global is validated in face interpolation and face shape modification experiments (see Sec. 4.3.1).
3.2.2 TDD-Regional Model
We integrate the above TDD-Global in our pipeline for automatic facial reconstruction. In addition, experts and investigators often require a manual tool to support local fine-tuning. The TDD-Global model alone does not allow for precise modification of specific facial landmarks without impacting other regions. On the other hand, modifying single tissue depth alone may result in a non-realistic face shape, because the tissue depths in a local region also adhere to a joint distribution. This feature is essential for detailed face editing. By enabling localized modification, the editing capability can be significantly enhanced.
To enable local face shape adjustments, we modify the TDD-Global model to account for the joint distribution of tissue depths in local regions. We follow a common facial region partitioning [34, 37] (forehead, cheeks, chin, middle region, and mouth) to divide facial landmarks into distinct groups $\mathcal{Y}_{l}\subset\mathcal{X}$, $k=1,\cdots,5$, as is visualized in Fig. 7(b). Then, by employing PCA to model the tissue depth distribution in each group individually, tissue depths can be adjusted locally within each subgroup.
According to the facial anatomy study [34, 37], facial musculature and neurovascular structures are located throughout faces and can be divided into several regions, each responding to different movements and expressions. Our TDD-Regional model can work effectively because it supports fine-tuning local face regions by adjusting regional tissue depths in accordance with relevant anatomical structures. We verified our TDD-Regional model by extensive experiments, as shown in Sec. 4.3.2 (e.g., Fig. 7) and the Supplementary video/document.
3.3 Initial Face Generation
When reconstructing a 3D face from a skull, in addition to geometric constraints, the biological profile derived from the skull can also aid in the reconstruction process. The biological attributes often include age, gender, and ancestry, which determine the basic characteristics of the generated face. In this first stage, we generate initial faces following the characteristics of the population group specified by these attributes.
We adopt Stable Diffusion XL [4] to generate high-resolution 2D portrait images from text prompts. We design four main dimensions of facial attributes: gender, ancestry, age, and face shape to generate various faces covering different populations. We also add additional prompts to ensure the diffusion model generates appropriate front faces.
We lift the synthesized portrait images to 3D faces by using DECA [15]. It encodes the image to latent space, consisting of shape, expression, pose, and albedo codes, then decodes the latent code to a 3D face with the FLAME representation. We transfer the texture and geometry details from the image to the 3D mesh. The generated faces conform to the biological profile and will be adapted to the final face by the geometric constraints of facial landmarks.
3.4 Anatomy-guided Face Adaptation
Given an initial 3D face model $F_{in}$, and a set of facial landmark constraints, $\mathcal{P}^{f}=\{\mathbf{p}_{i}^{f}\}$ recommended by our TDD model, we design an anatomy-guided face adaptation model to modify the initial face, ensuring the revised face $F_{out}$ aligns with all facial landmarks. Since our face models $\{F\}$ are generated by DECA, they have corresponding latent codes $\{f\}$, and can be decoded into FLAME models $\{G_{F}\}$; thus, all generated faces share consistent mesh connectivity, $F=G_{F}(f)$. We manually labeled the corresponding $n$ vertex indices on one face as facial landmarks, denoted as $\mathcal{Q}^{f}=\{\mathbf{q}^{f}_{i},i=1,\cdots,n\}$. These vertex indices remain consistent across all the generated faces.
Firstly, we estimate a transformation $H$ to make a coarse registration between $\mathcal{Q}^{f}_{in}$ of initial face $F_{in}$ and $\mathcal{P}^{f}$. We solve a Procrustes optimization to estimate the scaling, rotation, and translation parameters of $H$. With this linear transformation, we transform $\mathcal{P}^{f}$ to $\mathcal{\tilde{P}}^{f}$ as the roughly aligned landmark constraints. Secondly, we optimize the latent code $f$ by minimizing the objective function $\mathcal{L}$,
[…]
$\mathcal{L}$ $=\alpha_{0}\mathcal{L}_{lmk}+\alpha_{1}\mathcal{L}_{proj}+\alpha_{2}\mathcal{L}_{sym}.$ (2)
We design three losses in the objective function and will detail each of them below.
Landmark loss.
The deviation between the aligned landmark constraints $\mathcal{\tilde{P}}^{f}$ and the landmarks from the generated face should be minimized:
$\mathcal{L}_{lmk}=\|\mathcal{Q}^{f}-\mathcal{\tilde{P}}^{f}\|_{2}^{2}.$ (3)
Projection loss.
Since the corresponding landmarks on face $\mathcal{Q}^{f}$ are manually labeled, they are inevitably mismatched to the target facial landmarks $\mathcal{\tilde{P}}^{f}$. In this sense, $\mathcal{L}_{lmk}$ can be regarded as a regularization term. To improve the registration accuracy from the face geometry to the target landmarks, we project the target facial landmark constraints $\mathcal{\tilde{P}}^{f}$ to the reconstructed face as $\mathcal{\tilde{P}}^{pf}$. Then we minimize the Euclidean distance between $\mathcal{\tilde{P}}^{f}$ and $\mathcal{\tilde{P}}^{pf}$ as a data term to constrain the face shape,
$\mathcal{L}_{proj}=\sum_{p\in\mathcal{\tilde{P}}^{pf}}\underset{v\in F}{\min}\|v-p\|_{2}^{2},$ (4)
where $v\in F$ is vertex of the reconstructed face $F$.
Symmetry loss.
Facial symmetry plays an important role in authentic face synthesis that people prefer symmetry faces [17, 23]. Hence, we propose a symmetry loss term to encourage the symmetry of the synthesized face. We first split the facial landmarks $\mathcal{Q}^{f}$ into three groups, $\{\mathcal{Q}^{f}_{L},\mathcal{Q}^{f}_{M},\mathcal{Q}^{f}_{R}\}$, representing on the left, mid, right sides. The mid-plane $\Pi$ is perpendicular to the line through the eyes, which crosses the tip of the nose and the chin, which is computed by minimizing the least square error of $\mathcal{Q}^{f}_{M}$.
For any landmark on the left side $\mathbf{q}^{f}_{i}\in\mathcal{Q}^{f}_{L}$, we denote $\mathbf{q}^{pf}_{i}\in\mathcal{Q}^{f}_{R}$ as the corresponding landmark on the right side. We also denote the reflected point across the plane $\Pi$ as $\mathbf{q}^{rf}_{i}$. The symmetry loss function is defined as the Euclidean distance between the actual midpoint $\mathbf{q}^{M}$ and the expected midpoint $\mathbf{q}^{M^{\prime}}$,
$\mathbf{q}^{M}=\big{(}\mathbf{q}^{f}_{i}+$ $\mathbf{q}^{pf}_{i}\big{)}/2,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathbf{q}^{M^{\prime}}=(\mathbf{q}^{f}_{i}+\mathbf{q}^{rf}_{i})/2,$ (5)
$\mathcal{L}_{sym}$ $=\sum_{\mathbf{q}^{f}_{i}\in\mathcal{Q}^{f}_{L}}\|\mathbf{q}^{M}-\mathbf{q}^{M^{\prime}}\|_{2}^{2}.$ (6)
When the face latent code $f$ is optimized, it outputs the geometry of the 3D face $F_{out}$ that conforms to the landmarks and skull constraints. We transfer the appearance texture and geometry details previously extracted from the 2D portrait to $F_{out}$ as our facial reconstruction.
Figure 4: Examples of the synthesized faces by different attributes.
4 Experiment
4.1 Dataset and Error Metric
We adopted a CT-scan dataset, named Skull100, comprising 100 pairs of 3D skulls and faces. These pairs were reconstructed (using [31]) from whole-head CT scans. Due to the privacy issue, this dataset doesn’t collect face images or appearance data. Details of dataset collection are explained in the supplementary. Skull100 is used as the ground truth to evaluate faces reconstructed from their corresponding skulls. We divide the dataset into two subsets: 50 pairs for training and 50 for testing. This large test set size is chosen to ensure the statistical significance of our evaluation results. We conducted the K-Fold cross-validation to evaluate the distribution bias (Sec. 4.4.2).
We adopted the normalized mean error metric $E$ [29] to evaluate the geometric deviation of the reconstructed face $F$ and the ground truth $G$, where $F$ and $G$ are represented as triangular meshes. For each vertex $v$ in the ground truth face $G$, we find the closest vertex $u$ from the reconstructed face $F$ and sum the Euclidean distance as the reconstruction error,
$E=\frac{1}{d|G|}\sum_{v\in G}\leavevmode\nobreak\ \underset{u\in F}{\min}\leavevmode\nobreak\ \|v-u\|_{2},$ (7)
where $d$ is the distance between the two ears of the ground truth face, and $|G|$ is the total number of vertices in the ground truth face. Metrics Mean, Max, and Std are the average, maximum, and standard deviation of reconstruction errors $E$ of all test skulls.
4.2 Initial Face Generation
Since facial reconstruction should align with the biological profile of the skull, we design the attributes in the text prompts and leverage the Stable Diffusion XL to generate 2D portraits. We show the examples of the generated 2D face in Fig. 4. The complete text prompt and portrait attributes are reported in Supplementary. We will release this synthesized face dataset with attributes.
4.3 Evaluation of Tissue Depth Distribution
4.3.1 TDD-Global Model
The TDD-Global model allows users to use only one parameter $C$ to sample various valid combinations of tissue depths. The distribution and parameter choice of PCA decomposition are analyzed in the Supp.
In Fig. 5, we enumerated the TDD-Global model to sample different combinations of tissue depths from short to long. The 3D faces fit these tissue depths well and the face shape becomes fatter from left to right. Existing facial reconstruction methods [22, 56] only forcibly memorize the distribution of skull and face. Instead, our TDD-Global models and discovers the intrinsic pattern of tissue depths and human face shapes. It allows researchers to efficiently try different tissue depths and examine a variety of facial reconstruction possibilities.
Figure 5: TDD-Global model can explore faces with various shapes. From left to right, we iteratively sampled values of distribution $C$, and the corresponding tissue depths increased from short to long. The adapted faces that fit the tissue depth become fatter.
Figure 6: Validation of the face shape modification by TDD-Global model celebrities Christian Bale, Adele Adkins, and Ling Jia. We started from a 2D image (a) and used DECA to create its initial 3D face (b). We estimated proxy skull landmarks and adjusted the tissue depths upon them to tune the face shape to multiple roundness level (c). Face pose in (d) was intentionally adjusted for better visual comparisons with (e). (e) show the picture of the same person in a different face shape.
In the second experiment, we modified the face shapes of celebrities to show the ability of 3D face adjustment by the tissue depth model. Fig. 6 shows the results of three celebrities: Christian Bale, Adele Adkins, and Ling Jia. The initial 3D face of a celebrity was generated by the DECA model. We then used tissue depths of the pretrained TDD-Global model to estimate proxy skull landmarks (Sec. 3.2). We adjust $C$ to sample various valid combinations of tissue depths (e.g. from thin to fat) and then compute the target facial landmarks. Finally, the proposed anatomy-guided face adaptation is applied to adapt the initial 3D face to the new facial landmarks. The modified face is compared with the picture of the same person in a different face shape.
Figure 7: Adjusting local face shapes using the TDD-regional model. Tissue depths increase on the nose (from left to right) and cheek (from top to bottom), separately. These modified tissue depths fine-tune facial landmarks, thereby adjusting the face shapes.
4.3.2 TDD-Regional Model
We decoupled facial landmarks into distinct regions and applied separate PCA models to model the tissue depth distribution in each region, resulting in the TDD-Regional Model. This model offers explicit control for manipulating face shapes locally. As illustrated in Fig. 7, variations in the landmarks of the nose and cheek regions demonstrate the model’s capability. The nose size visibly increases from left to right, while the face shape appears fuller from top to bottom. We refer to the supplementary video for the animation of face fine-tuning.
Figure 8: Validating the influence of initial faces on the final results. (a) are the query skulls and facial landmarks. (b) is the synthesized portrait image. (c) are the initial 3D faces with texture and the reconstruction errors. (d) are our facial reconstructions, which and fit the landmark constraints well. The final reconstruction errors are similar: 0.899% (top) and 0.925% (bottom). They are both close to the GT face (e).
Figure 9: Qualitative comparisons with existing methods. (a) is the input skull. (b-e) show the results of HF-GGR [22], Pix2Pix [20], CycleGAN [58], and CGAN-CFG [56]. (f) show our reconstructed face with texture and geometry. (g) is the CT-scanned GT.
Figure 10: Facial reconstruction demo on historical figure: Robert the Bruce, was King of Scots from 1306 to his death in 1329. (a) is the 3D scan of the skull. (b) show the facial reconstruction by craniofacial experts Wilkinson et al. [55]. (c) is our synthesized portrait. We manually labeled the skull landmarks and used the pretrained TDD-Global to estimate the facial landmarks. (d) show the 3D reconstructed face that fits the facial landmarks. We adjusted the head pose for better visual comparison with reference (e).
4.4 Evaluation of Skull-to-Face Reconstruction
4.4.1 Influence of Initial Faces on Final Results
We chose the average tissue depths to compute the extended landmarks as facial constraints. Fig. 8 shows that with face refinement, our face adaptation model can effectively modify the face shape to align with the facial landmarks. With facial landmark constraints, the two initial faces, after adaptation, converge to a similar face geometry. Their geometries still exhibit subtle differences inherited from the initial faces (see error maps in (d)).
We conducted a quantitative experiment where we randomly synthesized four portraits for each skull and measured the final reconstruction errors. The reconstruction errors $E$ are similar for different initials: 1.756, 1.673, 1.724, and 1.755. For each skull, the overall face geometries amongst the final faces (from different initial faces) are similar. Subtle differences are noticed which are inherited from the initial faces (e.g. dimples, wrinkles, and lip thick).
We also conducted a user study to assess the convergence quality of the final facial reconstruction results from different initial faces. For each question, four initial faces are randomly synthesized. Three of them are used to reconstruct faces from Skull A, and the other one is used to reconstruct a face from Skull B. Users are required to identify the face geometries amongst the reconstructed faces and select the unique face that is reconstructed from Skull B. The detailed explanation and results of the user study are included in the Supplementary. The high average correctness of this user study indicates that our model can adapt different initial faces and converge to a face shape space that is related to the query skull, which is distinct from the reconstructed faces that are from the other skulls.
Table 1: Quantitative comparisons of facial reconstructions with existing methods. The distance between two ears of the groundtruth skin is normalized to 1.
Method Mean (%$\downarrow$ Max (%$\downarrow$ Std. (%$\downarrow$
HF-GGR [22] 2.198 4.755 0.841
Pix2Pix [21] 2.651 4.859 0.928
CycleGAN [59] 2.954 4.833 0.859
CGAN-CFG [56] 4.002 7.402 1.192
Ours (avg. tissue) 1.732 1.775 0.039
Ours (best tissue) 1.526 1.571 0.035
Table 2: Reconstruction errors before vs after adaptation and loss functions. Ablation Study Mean Recon. Mean Recon. (%) (millimeter) 1. Before adaptation 2.631 5.262   (from DECA) 2. w/ $\mathcal{L}_{proj}$ 2.273 4.547 3. w/ $\mathcal{L}_{lmk}$ 1.836 3.672 4. w/ $\mathcal{L}_{lmk}+\mathcal{L}_{proj}$ 1.758 3.576 5. After adaptation 1.732 3.464 ($\mathcal{L}_{lmk}+\mathcal{L}_{proj}+\mathcal{L}_{sym}$)
4.4.2 Comparisons with SOTA Facial Reconstruction Methods
Quantitative side-by-side comparisons are challenging due to the high cost and privacy concerns associated with joint skull-face CT scans, compounded by the absence of public benchmark data. Additionally, the source code for existing methods is not publicly accessible. Despite these limitations, we endeavored to conduct a fair comparison. CGAN-CFG [56], a state-of-the-art DNN-based facial reconstruction method, was tested using our training and testing dataset, with reconstruction results provided by the original authors. We also re-implemented CycleGAN [58] and Pix2Pix [20] following the guidelines in [56]. We chose HF-GGR [22] as the representative traditional PCA method and also re-implemented their model.
Tab. 4.4.1 shows the quantitative comparisons of facial reconstructions. Our method used synthesized portraits as initial faces. Ours (avg. tissue) used the average tissue depths from TDD-Global to compute the facial landmarks and then adapted the initial faces to the final results. Ours (best tissue) utilized three types of tissue depths thin, normal, and fat from TDD-Global as shown in Fig. 3. We selected the smallest reconstruction result as the final face. Our method outperforms existing methods in all metrics, largely due to the effective adaptation of our loss function to geometric constraints. The CT scan resolution is 0.25mm/pixel. The average of the reconstruction error is 3.464 millimeters.
Fig. 9 show qualitative comparisons. CGAN-CFG [56] tends to overfit the training set, leading to worse results on the test set. Similarly, Pix2Pix [20] and CycleGAN [58] often produce unrecognizable faces as they rely on the availability of large training datasets. HF-GGR [22] is a PCA-based model and typically generates average faces. More results are in the Supplementary.
To evaluate the distribution bias of training/test set, we used 5-fold cross-validation to split the dataset into training (80%) and testing (20%). The average reconstruction errors on the five testing sets are: 1.761%, 1.454%, 1.828%, 1.575%, and 1.665%. The results are consistent with Table 4.4.1.
4.4.3 Validation on Real-world Facial Reconstruction
We downloaded the 3D-scanned skull from the Anatomical Museums at the University of Edinburgh and used the proposed pipeline to reconstruct the faces for it. We used the text prompt to generate the initial face according to the historical background. Fig. 10 shows the comparisons between our results and the bust of Robert the Bruce at the National Wallace Monument. Wilkinson et al. [55] manually sculpted and painted the facial reconstruction. Although manual facial reconstruction could be subjective to different researchers [55], the reconstruction of Wilkinson and ours have similar chin, jawline, eyes, and face proportions, which are aligned to the latest depiction [55].
4.5 Ablation Study on Face Adaptation
We conducted an ablation study to evaluate the impact of various loss functions via three settings: (1) using only $\mathcal{L}_{lmk}$, (2) using $\mathcal{L}_{lmk}$ + $\mathcal{L}_{proj}$, and (3) our complete loss. Tab. 2 reports the average reconstruction errors. We computed the errors in the physical metric (millimeter) by the average face width (20cm). The projection loss $\mathcal{L}_{proj}$ aids in reducing the distance between the reconstructed face and landmark constraints. The symmetric loss $\mathcal{L}_{sym}$ enhances the face realism by promoting symmetry. The second and third rows show the face adaptation results of the same skull, using two different initial faces.
5 Conclusions
We proposed a novel anatomy-guided 3D face reconstruction approach that can reconstruct faces from a given skull using both biological profiles of the skull and the statistical tissue depth information defined on physiological landmarks. Our pipeline includes three main designs, the biological profile-based 3D face initialization, the facial landmark prediction based on a pre-computed tissue depth distribution, and the facial geometry adaptation to the landmarks. These designs enable the effective generation of multiple realistic faces with varying facial features, skin textures, and adiposity levels. Our results demonstrate that the proposed pipeline can effectively generate faces that closely conform to the given skull. This work provides a new tool for law enforcement agencies to generate and explore possible faces from skeletal remains, which greatly facilitates forensic identification when other identifying information is unavailable.
Limitations.
A limitation of our current approach is the constrained complexity of the tissue depth model due to the limited size of the available skull-face dataset. The high cost and prohibitive nature of collecting a large dataset have restricted our ability to train more sophisticated models. We will explore potential synthetic data generation techniques to enhance the dataset size and diversity, thereby enabling the exploration of more complex models for facial reconstruction.
Appendix 0.A Implementation Details
0.A.1 Text Prompt
We aim to synthesize 2D portraits that are depicted from the frontal view without occlusion. Hence, we designed the following prompt template to synthesize 2D portraits as initial faces,
prompt: A photo of a human face, facing the camera, full head visible, sharp focus, fine art portrait photography, with {Age}, {Ancestry}, {Gender}, and {Face Shape}.
[…]
negative prompt: Hat, glasses, black and white, cartoon, anime.
To reconstruct the face of Robert the Bruce, King of Scotts, from the scanned skull. We used the following text prompt to generate the initial faces from his historical background,
prompt: A photo of a king’s face. A Scottish male with a mustache, long hair, crown, and royal attire. Looking at the camera. Realistic portrait.
Table 3: Attributes used for initial face generation to match the biological profile of the skull.
Attributes Values
Age 10 to 20, 20 to 30, 30 to 50, 50 to 70
Ancestry African, Asian, European, Latino
Gender Female, Male
Face Shape Normal, Fat, Thin
0.A.2 Network Weights and Hyper-parameters
We used PCA to fit the tissue depths of the training set. To provide local control, we further divided the face into five regions and used PCA to fit tissue depths in each region. We kept the principal component with the largest eigenvalue for both global control and local control. We used the pre-trained models and fixed their weights in our pipeline (Stable Diffusion XL, DECA encoder, and FLAME decoder).
In Eqn. 2, we set $\alpha_{0}=5$, $\alpha_{1}=1$, $\alpha_{2}=1$. We use AdamW [32] to optimize the face latent code $f$ with an initial learning rate of $10^{-2}$. The learning rate is multiplied by $0.2$ for every $200$ iterations, with $1000$ iterations in total.
0.A.3 Dataset Processing
The whole-head CT scans were obtained by using a clinical multi-slice CT scanner system which provides CT images with a resolution of $512\times 512$ (pixel size $0.5\times 0.5mm$, inter-slice thickness $0.75mm$). After extracting meshes from the CT data, the heads and their underlying skulls are normalized into a unified Frankfurt coordinate system [11] to eliminate the inconsistency in posture and scale.
0.A.4 Pipeline Running Time
We calculated the overall running time on a single NVIDIA 4090 GPU. The 2D portrait is synthesized by a pre-trained Stable Diffusion XL model (7s). The initial 3D face is reconstructed from the synthesized 2D image by a pre-trained DECA model (2s). Then we adapt the face to match the landmark constraints and use a back-propagation to optimize the face latent code (5s). The overall time consumption is around 14 seconds for facial reconstruction from a skull.
Figure 11: PCA decomposition of tissue depths in the training set, with the 1st component dominating the distribution with $51.9\%$ of the variance.
Figure 12: Ablation study on loss functions. Projection loss $\mathcal{L}_{proj}$ helps reduce face deviation from landmark constraints; symmetric loss $\mathcal{L}_{sym}$ enhances face realism.
Appendix 0.B Analysis on Tissue Depth Distribution
We used PCA to decompose the joint distribution of tissue depths from the training set of the skull-skin dataset. We observed that the first component dominates the variances of PCA results, as Fig. 11 shown.
Appendix 0.C More Analysis on Face Adaptation Module
0.C.1 Qualitative Results of Loss Functions
Fig. 12 shows the reconstruction results using different loss functions. The proposed projection loss $\mathcal{L}_{proj}$ and symmetric loss $\mathcal{L}_{sym}$ help enhance the face realism. A quantitative ablation study is reported in the main text.
Figure 13: Example questions in the user study. Faces A, B, and C were reconstructed from one skull, while Face D was reconstructed from the other skull. Faces are shuffled in the user study.
Figure 14: User study of convergence quality of final facial reconstruction from different initial faces. We collected 500 valid responses from 50 valid users. The result shows that our model can adapt different initial faces and converge to a face shape space that is related to the query skull, which is distinct from the reconstructed faces that are from the other skulls.
0.C.2 User Study: Convergence Quality of Final Facial Reconstruction from different initial faces
Fig. 13 shows two examples of user study questions. For each question, we randomly synthesized four faces from four different portraits and used the proposed Anatomy-guided Face Adaptation module to adapt to the query skull. Faces A, B, and C are final reconstructions of one skull. Face D is the final reconstruction of the other skull. Faces are shuffled in the user study. Users are asked to compare the reconstructed faces and select the unique one reconstructed from the other skull.
We collected 500 valid responses from 50 valid users. For each question, we count 1 score so the overall score is 10 for 10 questions. Fig. 14 shows the user study results, where the average correctness is $91.2\%$. The result of the user study demonstrates that (1) our model can adapt different initial faces and converge to face shapes that are closely related to the query skull, resulting in distinct reconstructions for different skulls. (2) The reconstructed faces from one skull are not identical, they still preserved subtle geometric details inherited from the initial faces.
Refer to caption Refer to caption
(a) Changes of Tissue Depths (b) Faces Adapted to the Facial Landmarks
Figure 15: Adjusting local face regions by TDD-Regional. Tissue depths increase on forehead (from left to right) cheek (from top to bottom). The resulting faces closely align with the corresponding landmarks. Rectangles highlight the affected regions.
Refer to caption Refer to caption
(a) Changes of Tissue Depths (b) Faces Adapted to the Facial Landmarks
Figure 16: Adjusting local face regions by TDD-Regional. Tissue depths increase on the mouth (from left to right) and cheek (from top to bottom). The resulting faces closely align with the corresponding landmarks. Rectangles highlight the affected regions.
Appendix 0.D Detailed Results of Regional Control
We project the tissue depths in each face region to the first component of the PCA decomposition. Then users can intuitively and conveniently control the face shape by tuning one parameter. We show several examples below to demonstrate how users can intuitively and conveniently control the face shapes. For each example, we randomly select two different regions to visualize the results when tuning multiple regions at the same time.
Forehead and cheek control.
Fig. 15 shows the adjustment of the forehead and cheek regions by using different combinations of tissue depths. Our face adaption module can consistently fit the face to the landmark constraints well.
Mouth and cheek control.
We further show the adjustment of the mouth and cheek regions by using different combinations of tissue depths in Fig. 16.
Figure 17: Qualitative comparisons with existing methods. (a) is the input skull. (b-e) show the results of HF-GGR [22], Pix2Pix [20], CycleGAN [58], and CGAN-CFG [56]. (f) show our reconstructed face with texture and geometry. (g) is the CT scanned groundtruth.
Appendix 0.E Extensive Comparisons with Existing Methods
Fig. 17 shows extensive comparisons with the existing skull-to-face reconstruction methods. HF-GGR [22] is a method that tends to produce average faces that lack details. CGAN-CFG [56] directly memorizes the relationships between skull and face in the training set, leading to difficulties in generalization to the test set. Pix2Pix [20] and CycleGAN [58] focus on learning image-to-image translation for skull and face images, often resulting in unrecognizable faces due to their requirement of larger training datasets.